Cube root of 8a raise to power -6
Answer by Boreal(15235)

I rewrite as {(8a)^(1/3)}^-6
That is (8a)^-2=1/(8a)^2=1/64a^2
or {2(a)^1/3}^(-6), which gives the same answer.